KrazyKenzie:

4.9x + 4.4 = 19.1 a 4 b 3 c 4.8 d 7.2

4evablood:

wht da question

KrazyKenzie:

solve the equation 4.9x + 4.4 = 19.1

KrazyKenzie:

i gotts submit 25 of these before midnight i gotta hurry

4evablood:

ok

4evablood:

its 3

KrazyKenzie:

THANKSSSSSSSSSSSs

ramen:

1. 4.9x+4.4=19.1 2. Subtract the 4.4 from the left side of the equation 4.9x=19.1 - 4.4 3. 4.9x = 14.7 4. Just divide 14.7/4.9 to find X.

ramen:

I agree with 3.

4evablood:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KrazyKenzie THANKSSSSSSSSSSSs \(\color{#0cbb34}{\text{End of Quote}}\) np